### v1

# Guiding Generative Models for Protein Design: Prompting, Steering and Aligning

Abstract
Generative artificial intelligence (AI) models learn probability distributions from data and produce novel samples that capture the salient properties of their training sets. Proteins are particularly attractive for such approaches given their abundant data and the versatility of their representations, ranging from sequences to structures and functions. This versatility has motivated the rapid development of generative models for protein design, enabling the generation of functional proteins and enzymes with unprecedented success. However, because these models mirror their training distribution, they tend to sample from its most probable modes, while low-probability regions, often encoding valuable properties, remain underexplored. To address this challenge, recent work has focused on guiding generative models to produce proteins with user-specified properties, even when such properties are rare or absent from the original training distribution. In this review, we survey and categorize recent advances in conditioning generative models for protein design. We distinguish approaches that modify model parameters, such as reinforcement learning or supervised fine-tuning, from those that keep the model fixed, including conditional generation, retrieval-augmented strategies, Bayesian guidance, and tailored sampling methods. Together, these developments are beginning to enable the steering of generative models toward proteins with desired, and often previously inaccessible, properties.
1 Introduction
Learning to generate samples from complex probability distributions lies at the core of modern generative modeling. In the context of proteins, the availability of large datasets has catalyzed the rapid development of powerful generative models for protein design (GMPDs). Among GMPDs, some of the most adopted architectures are diffusion models, which reconstruct atomic coordinates by reversing a noise-adding process (Fig. 1A)[1, 2, 3], and protein language models (pLMs), trained either on masked sequence reconstruction (e.g. ESM[4] or the MSA Transformer[5]) or on autoregressive next-token prediction (e.g. ProtGPT2[6], ProGen2-3[7], and Evo-1/2[8, 9]) (Fig. 1B). In recent years, such models have enabled remarkable achievements, including the design of binders [10], ligand-binding receptors [11], and de novo enzymes [12] on unprecedented timescales.
Figure 1: Generative modeling approaches and schematic illustrations of protein fitness landscapes. (A) Diffusion-based generative models. Protein structures are represented by the distribution $p(x)$ of atomic coordinates in three-dimensional space. During training, true atomic coordinates $X_{0}$ are progressively corrupted by Gaussian noise until reaching a normal distribution $\mathcal{N}(0,1)$. At inference, the model reverses this process, iteratively denoising random noise to reconstruct a protein structure. (B) Sequence-based generative models. Here, $p(x)$ denotes the distribution of amino acids across sequence positions. Training can proceed via masked-token prediction, in which the model infers the identity of masked residues, or via autoregressive next-token prediction, where sequences are generated one residue at a time. (C) Protein fitness landscape. Schematic illustration of the relationship between the data distribution $p(x)$ (blue points), typically learned in an unsupervised manner, and the conditional distribution $p(y|x)$ (red points), which focuses sampling on regions associated with high fitness. (D) Evolutionary accessibility. Evolution explores the fitness landscape through local, incremental mutations that can reach only contiguous high-fitness regions (solid paths), whereas transitions across fitness valleys (dashed paths) are inaccessible. In contrast, generative protein design models (GMPDs) can, in principle, traverse the landscape in a less constrained manner, directly sampling from otherwise evolutionarily inaccessible regions.
Trained in a large corpus of natural sequences, structures, and functional annotations [e.g., 13, 14, 7, 15], GMPDs can explore vast regions of the protein sequence–structure landscape [e.g., 7, 16, 17, 18]. However, protein engineering often targets exceptional properties that are rare or even disfavored by natural selection (such as extreme thermostability in mesophilic organisms). These desirable functional optima may correspond to isolated high-fitness peaks separated by deep valleys in the evolutionary landscape (Fig. 1C, D), representing trajectories that evolution is unlikely to traverse. Consequently, GMPDs trained solely on natural data tend to assign negligible probability to these regions, making them difficult to sample directly. Further biases arise from uneven phylogenetic representation and residual annotation errors in the underlying datasets [e.g., 19, 20, 21, 22], which may further distort the learned landscape. In this sense, recent work highlight that simply increasing model scale does not guarantee monotonic improvements in fitness prediction [23].
However, when appropriately guided, GMPDs have the potential to access low-probability yet functionally optimal regions of protein space. More formally, the protein design objective can be defined as generating protein sequences $x$ with desired properties $y$, where $y$ may represent attributes such as thermostability, catalytic efficiency, binding specificity, or other design goals. By contrast, the primary training objective of a GMPD is to model the distribution of natural proteins, $p(x)$. To bridge this gap, recent approaches aim to model the conditional distribution $p(x|y)$, thereby biasing generation toward sequences with the desired attributes (Fig. 1C). In many cases, this objective can also be viewed as an optimization problem, where the goal is to find $x$ that maximizes $p(x|y)$. Several recent techniques can be conceptually unified under this $p(x|y)$ framework, showing strong success in aligning GMPDs with specific design objectives. In this review, we categorize these strategies into two broad classes: (i) train-time methods, which modify the model parameters so that the learned distribution $p(x)$ approximates $p(x|y)$, and (ii) inference-time control, which guides generation toward target properties without altering the model’s underlying weights.
1.1 Train-time methods for controlled Generation
Train-time methods modify the underlying probability distribution by directly updating GMPDs’ parameters. The most straightforward approach in this category is Supervised Fine-Tuning (SFT). In SFT, a pre-trained GMPD is further optimized under the same objective used during pre-training, but on a curated dataset of high-quality examples. For instance, a pre-trained protein language model (pLM) can be fine-tuned on a carefully assembled dataset from a target enzyme family to generate novel and distant members of that family. This process adapts the model’s parameters to a specific domain, effectively shifting its generative prior to align with the target data. SFT has achieved notable success, including the generation of enzymes [24, 25], gene editors[26, 27], or bacteriophages [28].
Method Description Examples
Train-time methods Supervised finetuning (SFT) Fit the model to well curated data, shifting the learned distribution towards the data.
Reinforcement Learning Aligns the model on feedback data over model’s outputs. Methods include preference- or reward-driven learning. PPO, GRPO, DPO
Inference-time control Prompt & context programming Generation guided by structuring the input prompt with explicit instructions or templates (e.g., for specific positions or motifs). Masked pLM, EvoDiff, BoltzGen
Retrieval-Augmented Generation (RAG) Enhances generation by dynamically incorporating external knowledge retrieved from a large corpus. RAG pipelines
Output-dependent guidance Gradient-based guided generation in sequence space based on the inference output. PPLM, ColabDesign, BoltzDesign, BindCraft
Activation steering Direct manipulation of hidden states (e.g., residual stream) to promote or suppress attributes without parameter updates. Sparse Autoencoders (SAEs), steering vectors
Bayesian guidance Re-weights the probability distribution using Bayes’ theorem: $\tilde{p}(y\mid x)\propto p_{\theta}(y\mid x)\exp(\lambda s(y,x))$. Bayesian scoring functions
Sampling controls Alters sampling strategy (temperature, top-k, top-p, or more advanced searches like beam search and MCTS) to influence randomness and diversity. MCTS, temperature scaling, top-k, top-p
Table 1: Overview of Train-time methods and Inference-time Control Methods with Examples. pLM: protein language model. MCTS: Monte Carlo Tree Search.
While SFT effectively specializes GMPDs toward generating samples representative of a particular dataset, it does not provide the model with the ability to discriminate by data quality, i.e., to differentiate among varying degrees of a desired property. Similar limitations have been observed in Natural Language Processing (NLP), where SFT alone often leads to suboptimal alignment with user intent and [29], in some cases, catastrophic forgetting [30]. Consequently, SFT is now commonly combined with reinforcement learning (RL) to achieve finer control over model behavior [e.g 31].
In RL, a model learns to make optimal decisions by interacting with an environment through trial and error, receiving feedback in the form of rewards or penalties to maximize its cumulative reward over time. Unlike SFT, in RL the model is not provided with explicit examples of the desired outputs. Instead, the GMPD must infer and explore autonomously, potentially uncovering novel solutions that might not have been anticipated. More technically, a pre-trained model is treated as a policy $\pi_{\theta}(x)$ and updated to maximize a scalar reward -or equivalently, the probability over preferences- while constraining excessive deviation from its pre-trained distribution. Similarly to SFT, this process enables a transition from the unconditional distribution $p(x)$ toward the desired conditional distribution $p(x|y)$. Today, RL is central to the alignment of large language models (LLMs) and has driven remarkable advances across diverse fields, from autonomous driving to game playing.
Several RL techniques have been implemented over the years. REINVENT (2017) [32] provides an early attempt, in the molecular field, to leverage vanilla RL policy-gradient methods (i.e., REINFORCE) to move from a broad distribution $p(x)$ toward a desired conditional distribution $p(x|y)$. REINVENT re-frames the popular REINFORCE update as an "augmented-likelihood objective". This objective allows to increase a property $y$’s score, while retaining probability mass near the original distribution $p(x)$, allowing the steering samples into high-scoring regions without sacrificing realism. Shortly after, Proximal Policy Approximation (PPO, 2017) [33] was introduced, providing a key improvement over vanilla policy-gradient methods; its clipped surrogate objective approximates a trust region, guaranteeing more stable updates.
Figure 2: Different stages where intervention can occur.
PPO became the basis for one of the two main families of RL from Human Feedback (RLFH, 2022) approaches, allowing to align LLM to human preferences. In the first approach, a reward model is first trained on (human) ranked responses (i.e., preferences) for desired properties $y$ (e.g., clarity of text, non-offensive text). Subsequently, the reward model is used within the PPO update to align a pre-trained LLM, encoding $p(x)$, towards the properties $y$. Similarly to REINVENT, RLHF augments the reward with a penalty for drifting too far from the original distribution $p(x)$. The more recent GRPO algorithm [34] (2024) computes the rewards for groups of LLM’s responses, augmenting the PPO update with a reward baseline (i.e., computed across the group), without needing to train any additional value model. Direct Preference Optimization [35] (DPO, 2023) introduced the second family of RLHF approaches. DPO exploits the same preference signal as PPO-based approaches, but recasts alignment as a supervised objective on the log-probability differences between ranked (human) preferences. In practice, this dispenses the need to train an explicit reward model, provides a simpler and, potentially, more stable route toward preference alignment.
These techniques have quickly met considerable success in the protein research realm, as summarized in table 2. PPO-style functional alignment has been used to update pLMs with experimental measurements, e.g., RL from eXperimental Feedback (RLXF) aligns an ESM-based generator toward brighter CreiLOV variants while constraining drift from the original distribution[36]. Model-based RL in the AlphaZero style has been applied to backbone construction, where Monte-Carlo Tree Search guided by a policy-value network outperforms plain tree search on top-down design tasks[37]. Preference-based tuning with DPO has been used to bias structure-conditioned models toward stability (preferring stabilizing over destabilizing sequences given a target backbone) and to reduce MHC-I epitope load while maintaining the fold[38, 39]. In parallel, mutation-policy RL frameworks propose on-policy sequence edits under specific oracles[36], and ProtRL extends the broad application of DPO and GRPO to protein engineering with pLMs, a framework that led to the design of low-nanomolar EGFR inhibitors[29].
Method Core objective Protein-design uses (2022-2025)
REINFORCE
(Vanilla Policy Gradient, 1992 [40])
Learns a policy such that actions with higher returns have higher likelihood of being sampled. AB-gen[41]
PPO
(Proximal Policy Optimization, 2017[42])
Plain policy gradients such as REINFORCE are quite noisy and unstable. PPO clips the model’s gradient update to not deviate too much from original model distribution. RLXF[36][43]
AlphaZero MCTS
(Monte Carlo Tree Search with policy–value networks, 2018 [44])
A neural networks with two heads is used to predict the most promising actions as well as the final rewards given a current position. This network is used to guide a MCTS to explore the most promising actions efficiently. EvoPlay[45], HighPlay[46]
DPO
(Direct Preference Optimization, [35])
Learns directly from ranked preference data, without requiring an explicit reward model. ProteinDPO[38], ProtRL[29] and Park et al.[47]
GRPO
(Group Relative Policy Optimization, [48])
For each prompt it samples a group of G candidate outputs, scores them with a reward, and uses relative (within-group) advantages to update the policy with a PPO-style clipped loss, without the need of a value model. ProtRL[29], ProteinZero [43]
Table 2: Representative reinforcement learning frameworks applied to protein sequence design. Summary of selected reinforcement learning (RL)–based approaches illustrating how core objectives have been adapted for protein or molecular design tasks between 2022 and 2025. The examples listed here are not exhaustive and focus on policy based RL, but highlight major methodological directions in the recent literature. For more details see [49], and Supplementary.
1.2 Inference-time control
In contrast to train-time methods, inference-time control methods guide GMPDs without updating their parameters, by directly influencing the models at the time of generation. These approaches assume that a pretrained GMPD already captures a rich representation of protein sequence-structure relationships, and that new behaviors can be elicited by steering the model’s inference process rather than retraining it. Such methods offer advantages in efficiency and flexibility, as the same model can be repurposed for diverse design tasks, enabling rapid exploration of new hypotheses while avoiding costly retraining.
Interventions can be applied at distinct stages of the generative process, from input to final sampling. Figure 2 summarizes these methodologies, categorized according to the site of intervention. Prompt and context engineering refers to shaping model behavior by modifying its inputs or conditioning information. pLMs can be trained with various control tags, such as Enzyme Commission numbers (ZymCTRL, [24]), Uniprot functional keywords (Progen, [25]), taxonomy (Evo1/2, [8, 9]), or combinations thereof [50]. Upon prompting these labels, the model specifically generates sequences for those cases. In other cases, specific residues may be masked and the model asked to reconstruct them, thereby exploring local sequence neighborhoods, or the model can be conditioned on predefined motifs and catalytic residues to guide functional design [51, 52, 53, 54, 10, 55]. Related to this idea, retrieval-augmented generation (RAG) has been introduced to dynamically incorporate external knowledge into the GMPD’s context[56, 57, 58]. By drawing upon semantically related examples from large databases, RAG enables more informed sampling and can enrich the design process with functional or structural priors not explicitly encoded in the pretrained model. A recent example is Protriever[57], which introduces a retrieval-augmented protein language model that jointly learns to retrieve homologous sequences and model their fitness, integrating evolutionary context at inference time without explicit structural supervision. Interventions can be applied at distinct stages of the generative process, from input to final sampling. Figure 2 summarizes these methodologies, categorized according to the site of intervention. Prompt and context engineering refers to shaping model behavior by modifying its inputs or conditioning information. pLMs can be trained with various control tags, such as Enzyme Commission numbers (ZymCTRL, [24]), Uniprot functional keywords (Progen1, [25]), taxonomy (Evo1/2, [8, 9]), or combinations thereof [50]. Upon prompting these labels, the model specifically generates sequences for those cases. In other cases, specific residues may be masked and the model asked to reconstruct them, thereby exploring local sequence neighborhoods[59], or the model can be conditioned on predefined motifs and catalytic residues to guide functional design [51, 60, 53]. Other approaches act directly on the latent encodings of protein sequences in a feedback-optimization loop, exploring regions of protein-structure space that satisfy user-defined design objectives such as ColabDesign, BindCraft and following methods [10, 54]. Alternatively, architectures such as BoltzGen[3] implement a conditional generative diffusion model with continuous guidance, in which encoded design conditions—such as binding-site specifications or structural constraints—are propagated throughout the denoising process to steer generation toward conformations consistent with the imposed design criteria [61] .
With the aim to dynamically inject additional knowledge into the GMPD’s context, Retrieval-augmented generation (RAG) has been successfully applied in different cases[56, 57, 58]. By drawing upon semantically related examples from large databases, RAG enables more informed sampling and can enrich the design process with functional or structural priors not explicitly encoded in the pretrained model. A recent example is Protriever[57], which introduces a retrieval-augmented protein language model that jointly learns to retrieve homologous sequences and model their fitness, integrating evolutionary context at inference time without explicit structural supervision.
A more surgical form of intervention acts directly within the hidden states of the network. Activation steering manipulates internal representations, often within the residual stream, by injecting vectors that correspond to interpretable latent directions. Sparse autoencoders (SAEs) have been used to identify such interpretable features in protein language models, revealing latent dimensions correlated with properties like enzymatic activity, hydrophobicity, or thermostability [62, 63]. For example, Parsan et al. used SAE-derived features to bias structure predictions in ESMFold toward more hydrophobic conformations through feature steering[64], while Boxò et al. leveraged activity-associated features to steer ZymCTRL toward more active $\alpha$-amylases[65]. This approach provides a surgical way of shaping the trajectory of the model’s activations without altering inputs or outputs explicitly.
Control can also be applied at the level of output probabilities. Bayesian guidance reweighs the probability distribution encoded by GMPDs using Bayesian principles, effectively combining the model’s prior with external evidence or predictive scores. Such strategies have been applied in protein design, where sequence likelihoods are updated according to functional predictors or activity [66].
Finally, Sampling Controls manipulate the stochasticity of the GMPD’s final output by, for example, manipulating "inference parameters" like temperature, top-k and top-p sampling, balancing diversity and fidelity, which is particularly important when sampling from vast sequence landscapes [67]. More advanced sampling techniques, such as beam search and Monte Carlo Tree Search (MCTS), allow to consider multiple GMPDs inference trajectories, selecting the optimal ones [e.g., 9, 68].
Recent theoretical work underscores the generality of these approaches: flow matching in discrete state spaces has been shown to be equivalent to masked language modeling, autoregressive generation, and diffusion. This unifying perspective positions inference-time control as a suite of architecture-agnostic, plug-and-play techniques that can be ported across model classes with minimal modification [66].
1.3 Conclusion and future prospects
Protein design operates at the intersection of evolutionary complexity and computational abstraction. Natural proteins are shaped by diverse and often competing forces: biochemical constraints, ecological pressures, and evolutionary contingencies; making it difficult to define a single, global “fitness vector”. Instead, the true fitness landscape is fluid, heterogeneous, and context-dependent. Combined with well-known dataset biases [e.g., see 19, 20, 22], these properties impose severe limitations on how much we can leverage the natural distribution of proteins to design and engineer proteins à la carte, with full control over the design process.
Here, we discussed two broad categories of methods, fine-tuning and inference-time control, which are arising as powerful tools to guide GMPDs towards desired regions of the learned protein distributions (e.g., those encoding target design properties), giving us tighter control over the protein design process.
However, a recurring limitation of both families is poor out-of-distribution (OOD) generalization. Inference-time control presupposes that GMPDs already encode rich structure over the relevant regions of sequence space; steering cannot along produce representations that the base model never learned. Fine tuning methods may allow for more flexibility, enabling to reshape the GMPD’s learned distribution by changing its parameters. However, many curated datasets used for supervised fine-tuning and in-silico scoring methods used for RL are based on the natural distribution of proteins, reflecting the same biases and evolutionary constraints we would like to move away from [e.g., 69, 18, 60].
In this regard, including physics-based scoring methods, like RoseTTA fold [70] and FoldX [71], in the RL fine-tuning process may provide a promising avenue, enabling to move away from models that exclusively infer information from the natural distribution of proteins. [72].
Moreover, as GMPDs are pushed to explore protein regions distant from the natural distribution of proteins, experimental testing providing reliable validation is paramount. In this regard, lab-in-the-loop frameworks combining GMPDs with experimental validation are a very promising avenue. Advances in laboratory automation [73] and techniques[74, 75] are reducing experimental bottlenecks, enabling faster experimental testing and unprecedented scale of screening.
A complementary strategy is to improve the data distribution seen during pretraining and downstream optimization: diversifying sequence sampling across the tree of life [76]. Other approaches introduce inductive biases, leveraging evolutionary context provided by Multiple Sequence Alignment to provide more efficient training [77] or sampling[22].
The field of generative modeling for protein research is advancing at an unprecedented pace. A multitude of techniques for the guidance and control of GMPDs has emerged in the past two years, demonstrating promise for the targeted engineering of proteins across diverse applications. As the field progresses, a central challenge will be extending these models beyond the natural sequence distribution—ensuring reliable performance in out-of-distribution (OOD) regimes and paving the way toward truly generalizable, controllable, and creative protein design.
1.4 Declaration of competing interest
The authors declare no conflict of interest.
1.5 Acknowledgments
We thank Gerard Boxó and Alex Vicente and other members of the Ferruz lab for their thoughtful feedback. FS is supported by a fellowship from the “la Caixa” Foundation (ID 100010434, fellowship code DFI25-00620F). MG acknowledges funding from project CPP2022-009990 funded by MCIN/AEI/10.13039/501100011033 and by the European Union (FEDER, UE). NF acknowledges support from the Ramón y Cajal Fellowship (RYC2021-034367-I) funded by MICIU/AEI/10.13039/ 501100011033 and the European Union NextGeneration EU/PRTR, and from the European Union’s Horizon Europe programme under grant agreement No 101165231. We acknowledge support of the Spanish Ministry of Science and Innovation through the Centro de Excelencia Severo Ochoa (CEX2020-001049-S, MCIN/AEI /10.13039/ 501100011033), and the Generalitat de Catalunya through the CERCA programme. We are grateful to the CRG Core Technologies Programme for their support and assistance in this work. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.